Peter sold pumpkins from his farm. He sold Jumbo pumpkins for $9 each and sold Regular pumpkins for $4 each. Peter sold 80 pumpkins and collected $395. How many Jumbo pumpkins and Regular pumpkins did he sell?
Answer by stanbon(75887)

Peter sold pumpkins from his farm. He sold Jumbo pumpkins for $9 each and sold Regular pumpkins for $4 each. Peter sold 80 pumpkins and collected $395. How many Jumbo pumpkins and Regular pumpkins did he sell?
---------------------
Let # of jumbos be "x"; value = 9x dollars
--------------------
# of regulars is "80-x"; value = 4(80-x)= 320-4x dollars
----------------------
EQUATION:
value + value = 395 dollars
9x + 320-4x = 395
5x = 75
x = 15 (# of jumbos)
80-x = 65 (# of regulars)
==================
Cheers,
Stan H.